Please use the table to answer questions, SHOW WORK

Consider the following work breakdown structure for questions 46-50: Duration (Days) Most Start Finish Activity Node Node Optimistic Likely Pessimistic 50 31 70 41 35 59 34 82 41 58 32 30 5 46
What is the estimated expected (mean) time tor activity D? A 32 days B. 41 days C 44 days D. 42 days E 40 dayts 46 What is the estimated standard deviation for the critical path? A 6 days B. 5 days C. 49 days D. 3 days E. 2 days 47 48 What is the estimated expected (mean) time for project completion? A 70 days B 100 days C 106 days D 120 days E. 226 days What is the estimated slack time for activity D? A 0 days B. 10days C. 14 days D. 20 days E. 80 days 49 What is the probability that this project will be completed within 130 days? A 08413 В 09 C 09544 D 0.9772 50
Answer #1

A 50 2 5 (90 2 6 D (46 了

50 30 70 40 1 Activity Optimistic Most Likely Pessimistic Expected (Mean) time Variance 59 34 82 50 31 70 41 35 41 16 58 32 3

Expected mean time = (Optimistic + pessimistic + 4*most likely)/6

Variance = ((Pessimistic - Optimistic)/6)^2

46.

Correct answer is E.40 days

47.

Path available: A-C duration => 120 days

B-D-E duration => 106 days

Critical path is A-C
Variance of critical path = 9+16 = 25 days
Standard deviation of Critical path A-C is sqrt(25) = 5 days
So correct answer is B. 5 days

48.

Estimated time for project completion = Duration of critical path A-C =120 days

Correct answer is D. 120 days

49.

Earliest Start time of B = 0 days
Earliest Finish time of B = Earliest Start time of B + Duration of B = 0+30 = 30 days

Earliest start time of D = Earliest Finish time of B = 30 days
Earliest Finish time of D = Earliest Start time of D + Duration of D = 30+40 = 70 days

Late Finish of E = 120 days
Late Start of E = Late Finish of E - Duration of E = 120 - 36 = 84 days

Late Finish of D = Late Start of E = 84 days

Slack of D = Late Finish - Early Finish = 84 - 70 = 14 days

Correct answer is C. 14 days

50.

When x = 130

z = (130 - 120)/5 = 2

Probability of project completing within 130 days = P(Z<2) = 0.9772

So correct answer is D) 0.9772